The area of a rectangle wall of a barn is 144 square feet. Its length is 10 feet longer than the width. What is the length and width of the wall of barn?
Please help me!

Answer by Alan3354(69443)

L = W + 10
L*W = 144
W*(W+10) = 144
W^2 + 10W - 144 = 0
(W + 18)*(W - 8) = 0
W = 8
L = 18